OpenStudy (lexia99):

1. what is the sum or difference 2x^4-8x^4

OpenStudy (thefurball):

They are like terms, so you can just add/subtract their coefficients. Their coefficients are 2 and -8.

OpenStudy (lexia99):

okay I think i got it thanks

alones (alones):

2x4+−8x4 \(2x^{4}\) +- \(8x^{4}\) \(\ x^{4}\)+ \(\ x^{4}\)\(+−x^{4}\)\(+−x^{4}\)\(+−x^{4}\)\(+−x^{4}\)\(+−x^{4}\)\(+−x^{4}\)\(+−x^{4}\)\(+−x^{4}\)\(−x^{4}\)\(+−x^{4}\)\(+−x^{4}\)\(+−x^{4}\)\(+−x^{4}\)\(+−x^{4}\) \(−6x^{4}\)

alones (alones):

Loool sorry it took mee so loong xP

OpenStudy (thefurball):

Yeah I think @AloneS did it the easiest way

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\)\(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\) @lexia99

OpenStudy (lexia99):

its fine it helped @AloneS. i have another one 10y^10-14y^10. would it be -4y^20 ?

OpenStudy (thefurball):

Ya

OpenStudy (anonymous):

2x^4-8x^4=x^4(2-8)=-6x^4

OpenStudy (thefurball):

wait no

OpenStudy (thefurball):

the y is still to the 10th

OpenStudy (lexia99):

so the only thing I would be doing is adding or subtracting the 14 and the 10? so lost lol